# Deterministic and Statistical Analysis of the DoF of Continuous Linear Arrays in the Near Field

### v1

@@ IV Calculation of the DoF
The evaluation of the DoF consists of decomposing the spatial channel into a series of independent parallel sub-channels by determining the equivalent “singular values” through an eigenvalue decomposition of the Hermitian kernel of Green’s function. The resulting eigenfunctions constitute two complete sets of orthogonal basis functions, one associated with the Tx array and the other with the Rx array. The number of non-zero eigenvalues of Green’s function kernel, is defined as the number of DoF. However, the analytical solution to this problem for a generic geometric array setup is highly challenging. An alternative solution was proposed in [40] using the kernel functions. The kernel function for two points $\zeta,\zeta^{\prime}$ on the Rx array is given by
$K_{R}(\zeta,\zeta^{\prime})$ $=\int_{-l_{T}/2}^{l_{T}/2}G(\zeta,\eta)G^{*}(\zeta^{\prime},\eta)\rm{d}\eta$ (31)
$=\int_{-l_{T}/2}^{l_{T}/2}\frac{e^{-jkr}}{4\pi r}\frac{e^{jkr^{\prime}}}{4\pi r^{\prime}}\rm{d}\eta$
The distances $r$ and $r^{\prime}$ in the denominators can be approximated by the distance between the centers of the two arrays $d_{0}=\sqrt{x_{0}^{2}+y_{0}^{2}}$. This approximation is valid only for the amplitude and cannot be adopted for the phases in (31). The kernel function can be written with the help of the focusing function defined in (8), as follows:
$K_{R}(\zeta,\zeta^{\prime})$ $\approx\frac{1}{(4\pi d_{0})^{2}}\int_{-l_{T}/2}^{l_{T}/2}e^{-j\frac{2\pi}{\lambda}(\rho-\rho^{\prime})\eta}\rm{d}\eta$ (32)
$=\frac{l_{T}}{(4\pi d_{0})^{2}}{\rm{sinc}}\left(\frac{l_{T}}{\lambda}(\rho-\rho^{\prime})\right),$
where ${\rm{sinc}}(x)=\sin(\pi x)/\pi x$. This kernel is identical to the field distribution at the Rx array when the phase profile at the Tx array is set to focus towards $\zeta^{\prime}$
@@
In order to compute the number of DoF supported by the two arrays, we calculate the number of corresponding orthogonal steering functions that fit within the length of the Rx array. This is accomplished by setting a reference point $\zeta^{\prime}$ and finding the number and location of the points on the Rx array where the kernel function is zero. The usefulness of the ${\rm{sinc}}$ function in (32) is that one may easily calculate the points $\zeta$, where it is zero, i.e., the integer multiples of $\frac{l_{T}}{\lambda}(\rho-\rho^{\prime})$. It is worth noting that the proposed approach generalizes the work in [40] by allowing the Rx array: i) to rotate and ii) its center coordinate, $x_{0}$, to take negative values. Moreover, the Tx and Rx arrays radiate only in one of the two half-planes.
To this end, we set the reference point at the center of the Rx array, i.e., $\zeta^{\prime}=\zeta_{c}$, and solve the following equation for the points $\zeta_{m}$, which are measured with respect to the center $\zeta_{c}$ and fall within the array length, $-l_{R}/2\leq\zeta_{m}\leq l_{R}/2$:
@@
where the values of $a_{+}$ and $a_{-}$ are determined from the corresponding values of $\gamma_{+}$ and $\gamma_{-}$ that are obtained from (35) by setting $\zeta_{m_{+}}=\frac{l_{R}}{2}$ and $\zeta_{m_{-}}=\frac{-l_{R}}{2}$, respectively. In addition, $\rho_{c}$ and $\gamma_{c}$ are calculated by setting $\zeta_{m}=0$ in (35).
Therefore, the number of DoF is given by
$m=|m_{+}-m_{-}|.$ (39)
In the far-field setting, i.e., when the Tx array is in the far field of the Rx array and the transmitting and receiving sides of the two arrays are mutually visible, a constant $1$ is to be added to (39) to ensure consistency [29]. In Fig. 3, a typical calculation of the number of DoF is illustrated as a function of $\theta_{R}$, based on (39).
@@
IV-A A Geometric Setup of Interest
In this subsection, a practical geometric setup is considered where a small array is located in front of a large array, which resembles the placement of user equipment in front of a radio unit. This case study is of interest since, for a range of values of the variable $\theta_{T}$, it provides a paraxial setup, i.e., the midpoints of the two arrays are perfectly aligned along the same line and thus $y_{0}=0$, whereas for other values of $\theta_{T}$ the center of the Rx is shifted, leading to a non-paraxial case. This is due to the partial visibility of the Rx array and thus the need to calculate the DoF with respect to the new center. This setup will also be utilized in the next section to study the statistical behavior of the DoF. Moreover, assume that $x_{0}\geq L_{T}/2$, that is, Rx is in the right half-plane, and $\theta_{R}=\pi$. The selection of $L_{T}/2$ as a lower limit for the coordinate $x_{0}$ is made to avoid mutual crossing of the two arrays. Since the Tx array has a small size, in fact, the impact of the mutual crossing may reasonably be assumed to be negligible. The coordinate $x_{0}$ and the rotation angle of the TX surface determine the visibility condition, and consequently the number of DoF between the two surfaces. Therefore, there are four possible cases: i) the Rx array is not visible from the Tx and the DoF is 0, ii) the Rx array is partially visible from the Tx array and the endpoint $R^{+}$ is visible, iii) the Rx array is fully visible from the Tx, and iv) the Rx array is partially visible from the Tx and the endpoint $R^{-}$ is visible. The first case occurs for $-\pi<\theta_{T}\leq a_{+}^{max}-\frac{\pi}{2}$ and for $a_{-}^{max}+\frac{\pi}{2}\leq\theta_{T}\leq\pi$, where $a_{+}^{max}=\arctan\gamma_{+}=\arctan(-L_{R}/2x_{0})$ and $a_{-}^{max}=\arctan\gamma_{-}=\arctan(L_{R}/2x_{0})=-a_{+}^{max}$. Then,
@@ IV-A A Geometric Setup of Interest
$m$ $=\frac{L_{T}}{\lambda}[\sin\left(\theta_{T}-a_{+}\right)-\sin\left(\theta_{T}-a_{-}\right)]$ (43)
$\mathop{=}\limits^{a_{+}=-a_{-}}\frac{2L_{T}}{\lambda}\cos\theta_{T}\sin a_{-}.$
@@ V-A Stochastic Geometry Framework
$m=\frac{L_{T}}{\lambda}|\sin(\theta_{T}-a_{+})-\sin(\theta_{T}-a_{-})|=\frac{L_{T}}{\lambda}|\rho_{+}-\rho_{-}|.$ (45)
@@ V-C1 Partial Visibility of the Rx Array
for $m\in\Big{[}0,2C_{L_{T},\lambda}\frac{L_{R}^{2}}{L_{R}^{2}+4x_{0}^{2}}\Big{]}$.
Proof. Assume that $R^{+}$ is visible. Conditioned on $x_{0}$, one can set $f_{x_{0}}(x_{0})=\delta(x_{0})$ in (46), which provides the PDF $f_{\rho_{-}}(\rho)$. The limits of the range for $m$ are obtained with the aid of (56) shown in Appendix A. Therefore, $f_{m|x_{0},R^{+}}(m)=f_{m|R^{+}}(m)\Big{|}_{f_{x_{0}}(x_{0})=\delta(x_{0})}$ and similarly $f_{m|x_{0},R^{-}}(m)=f_{m|R^{-}}(m)\Big{|}_{f_{x_{0}}(x_{0})=\delta(x_{0})}$. Finally, $f_{m|x_{0},R^{+}}(m)$ and $f_{m|x_{0},R^{i}}(m)$ have the same statistical behavior. $\blacksquare$
@@ V-C2 Full Visibility of the Rx Array
Conditioned on $x_{0}$, and with full visibility of the Rx array, the conditional PDF $f_{m|x_{0},FV}(m)$ of the number of DoF is given by Lemma 3 as $f_{m|x_{0},FV}(m)=f_{m|x_{0}}(m)$.
Conditioned on $x_{0}$, the PDF of the number of DoF is given by the following theorem.
Theorem 4: Conditioned on $x_{0}$, the conditional PDF $f_{m|x_{0}}(m)$ of the number of DoF, $m$, is given by (54), shown at the bottom of the next page, where $\mathcal{V}_{R^{+}}=\mathcal{V}_{R^{-}}=\frac{\arctan(L_{R}/2x_{0})}{\pi}$ denotes the probability that $R^{+/-}$ is visible and $\mathcal{V}_{FV}=\frac{\pi-2\arctan(L_{R}/2x_{0})}{2\pi}$ denotes the probability that full visibility is achievable.

### v2

@@ IV Calculation of the DoF
Each array is considered as a continuum composed of an infinite number of infinitesimal antennas, each producing a spherical wave, i.e., it is a Huygens’ source. Although the transmitted field is a vector, it is approximated as a complex scalar to simplify the analysis. Correspondingly, in this work, Green’s function is not a tensor but a scalar function. Polarization effects may be incorporated by replacing every antenna point by three mutually perpendicular electric dipoles [49]. The evaluation of the DoF consists of decomposing the spatial channel into a series of independent parallel sub-channels by determining the equivalent “singular values” through an eigenvalue decomposition of the Hermitian kernel of Green’s function. The resulting eigenfunctions constitute two complete sets of orthogonal basis functions, one associated with the Tx array and the other with the Rx array. The number of non-zero eigenvalues of Green’s function kernel, is defined as the number of DoF. However, the analytical solution to this problem for a generic geometric setup is highly challenging. An alternative solution was proposed in [43] using the kernel functions. The kernel function for two points $\zeta,\zeta^{\prime}$ on the Rx array is given by
$K_{R}(\zeta,\zeta^{\prime})=\int_{-\frac{l_{T}}{2}}^{\frac{l_{T}}{2}}G(\zeta,\eta)G^{*}(\zeta^{\prime},\eta){\rm{d}\eta}=\int_{-\frac{l_{T}}{2}}^{\frac{l_{T}}{2}}\frac{e^{-jk(r-r^{\prime})}}{(4\pi)^{2}\,rr^{\prime}}\rm{d}\eta$ (33)
The distances $r$ and $r^{\prime}$ in the denominators can be approximated by the distance between the centers of the two arrays $d_{0}=\sqrt{x_{0}^{2}+y_{0}^{2}}$. As explained in [50], if the distance $d_{0}$ between the two arrays satisfies $d_{0}\geq 1.2\,(L_{T}+L_{R})$, the variations of the amplitude are negligible and thus the approximation $r\approx r^{\prime}\approx d_{0}$ for the amplitude is valid. This approximation cannot be adopted for the phases in (33). The kernel function can be written with the help of the focusing function defined in (10), as follows:
$K_{R}(\zeta,\zeta^{\prime})\approx\frac{1}{(4\pi d_{0})^{2}}\int_{-l_{T}/2}^{l_{T}/2}e^{-j\frac{2\pi}{\lambda}(\rho-\rho^{\prime})\eta}e^{-j\frac{2\pi}{\lambda}(\tilde{\rho}-\tilde{\rho}^{\prime})\eta^{2}}\rm{d}\eta$ (34)
This kernel is identical to the field distribution at the Rx array when the phase profile at the Tx array is set to focus towards $\zeta^{\prime}$
@@
Using [51, eq. 2.33.3] the kernel function in (34) is obtained in closed form in (36), shown at the bottom of the next page, where $\rm{Erfi}(\cdot)$ denotes the imaginary error function.
$K_{R}(\zeta,\zeta^{\prime})\approx\frac{1}{(4\pi d_{0})^{2}}\frac{(-1)^{1/4}e^{\frac{j(\rho-\rho^{\prime})^{2}k}{4(\tilde{\rho}-\tilde{\rho}^{\prime})}}\sqrt{\pi}\left({\rm{Erfi}}\left[\frac{(-1)^{3/4}\sqrt{k}\left[(\rho-\rho^{\prime})-(\tilde{\rho}-\tilde{\rho}^{\prime})l_{T}\right]}{2\sqrt{\tilde{\rho}-\tilde{\rho}^{\prime}}}\right]-{\rm{Erfi}}\left[\frac{(-1)^{3/4}\sqrt{k}\left[(\rho-\rho^{\prime})+(\tilde{\rho}-\tilde{\rho}^{\prime})l_{T}\right]}{2\sqrt{\tilde{\rho}-\tilde{\rho}^{\prime}}}\right]\right)}{2\sqrt{\tilde{\rho}-\tilde{\rho}^{\prime}}\sqrt{k}}$ (36)
In order to compute the number of DoF supported by the two arrays, we calculate the number of corresponding orthogonal focusing functions that fit within the length of the Rx array. This is accomplished by setting a reference point $\zeta^{\prime}$ and finding the number and location of the points on the Rx array where the kernel function is zero. The analytical computation of the number of zeros on the Rx array for the function (36) is cumbersome. Fortunately, hereafter, it is shown that an alternative way to compute the number of zeros is to approximate the kernel function with the linear term of the phase, as if the Rx array were positioned in the far field of the Tx array. Under this assumption, the kernel function is given by
$K_{R}^{ff}(\zeta,\zeta^{\prime})\approx\frac{l_{T}}{(4\pi d_{0})^{2}}{\rm{sinc}}\left(\frac{l_{T}}{\lambda}(\rho-\rho^{\prime})\right),$ (37)
where ${\rm{sinc}}(x)=\sin(\pi x)/\pi x$, and the superscript $(ff)$ denotes the far field approximation. The usefulness of the ${\rm{sinc}\left(\frac{l_{T}}{\lambda}(\rho-\rho^{\prime})\right)}$ function in (37) is that one may easily calculate the points $\zeta$, where it is zero, i.e., the integer multiples of $\frac{l_{T}}{\lambda}(\rho-\rho^{\prime})$. Indeed, the amplitude of the kernel function in (36) oscillates in the same manner as the absolute value of the approximated kernel function in (37) as a function of the difference $(\rho-\rho^{\prime})$. Furthermore, the number of minima of the initial kernel and the approximated one is identical. This is clearly depicted in Fig. 3 where the two functions are plotted for different cases and $\zeta^{\prime}=0$. The minima of the kernel functions occur at integer multiples of $\lambda/l_{T}$, where $l_{T}=L_{T}$ for the considered setups. It is noteworthy that when $(\rho-\rho^{\prime})\gg(\tilde{\rho}-\tilde{\rho}^{\prime})l_{T}$, as in Figs. 3(a), 3(b) and 3(d) the approximation is highly accurate as the arguments of the ${\rm{Erfi}}$ functions in (36) are nearly identical. Consequently, the minima of the function $|K_{R}(\zeta,\zeta^{\prime})|$ are congruent to those of the function $|K_{R}^{ff}(\zeta,\zeta^{\prime})|$ and gradually approach zero. Interestingly, when $(\rho-\rho^{\prime})>(\tilde{\rho}-\tilde{\rho}^{\prime})l_{T}$, as in Fig. 3(c), the number of minima remains the same, but their value are not necessarily zero for all of them. For the initial minima, i.e., for values of $\zeta$ close to the center of the Rx array, the deviation is greater, whereas for larger values of $\zeta$ it tends to zero. This behavior implies that some focusing functions are semi-orthogonal. However, since the effective length, $l_{T}$, of the Tx array plays a pivotal role on the ratio $[(\rho-\rho^{\prime})/(\tilde{\rho}-\tilde{\rho}^{\prime})l_{T}]$, the smaller the length, the smaller the deviation of the minima from the zero value.
(a) $\theta_{T}=0,\theta_{R}=\pi,x_{0}=10\,m,y_{0}=0\,m,L_{T}=0.2\,m,L_{R}=5\,m,f=30\,$GHz.
(b) $\theta_{T}=\pi/3,\theta_{R}=\pi,x_{0}=10\,m,y_{0}=0\,m,L_{T}=0.2\,m,L_{R}=5\,m,f=30\,$GHz.
(c) $\theta_{T}=\pi/3,\theta_{R}=-\pi/3,x_{0}=-5\,m,y_{0}=5\,m,L_{T}=1\,m,L_{R}=5\,m,f=30\,$GHz.
(d) $\theta_{T}=\pi/3,\theta_{R}=-\pi/3,x_{0}=-5\,m,y_{0}=5\,m,L_{T}=0.2\,m,L_{R}=5\,m,f=30\,$GHz.
Figure 3: Comparison of the amplitude of the kernel functions given by (37) and (36) for four distinct cases. The inset is a zoomed picture to highlight the deviation of the two functions.
It is worth noting that the proposed approach generalizes the work in [43] by extending the phase profile with a non-linear term and allowing the Rx array: i) to rotate and ii) its center coordinate, $x_{0}$, to take negative values. Moreover, the Tx and Rx arrays radiate only in one of the two half-planes. To this end, we set the reference point at the center of the Rx array, i.e., $\zeta^{\prime}=\zeta_{c}$, and solve the following equation for the points $\zeta_{m}$, which are measured with respect to the center $\zeta_{c}$ and fall within the array length, $-l_{R}/2\leq\zeta_{m}\leq l_{R}/2$:
@@
where the values of $a_{+}$ and $a_{-}$ are determined from the corresponding values of $\gamma_{+}$ and $\gamma_{-}$ that are obtained from (39) by setting $\zeta_{m_{+}}=\frac{l_{R}}{2}$ and $\zeta_{m_{-}}=\frac{-l_{R}}{2}$, respectively. In addition, $\rho_{c}$ and $\gamma_{c}$ are calculated by setting $\zeta_{m}=0$ in (39). Therefore, the number of DoF is given by
$m=|m_{+}-m_{-}|+1.$ (43)
The constant 1 is added to ensure that the reference position is also taken into account, i.e., the mode that corresponds to $(\rho-\rho^{\prime})=0$. In Fig. 4, a typical calculation of the number of DoF is illustrated as a function of $\theta_{R}$, based on (43).
IV-A Insights on the Calculation of the Number of DoF
The calculation of the number of DoF as outlined in the preceding section is explicitly dependent not only on the relative positions of the two arrays, as determined by the coordinates $(x_{0},y_{0})$, but also on their respective lengths $L_{T},L_{R}$, and on the rotation angles $\theta_{T}$ and $\theta_{R}$. In this section, a concise analysis of the dependency of the number of DoF on all the aforementioned variables is presented. This analysis aims to elucidate the practical implications of specific constraints on the number of DoF.
The initial observation is that the variables $x_{0}$ and $y_{0}$ need to be selected to ensure $d_{0}=\sqrt{x_{0}^{2}+y_{0}^{2}}>1.2(L_{T}+L_{R})$, for the approximation $r\approx r^{\prime}\approx d_{0}$ for the amplitude in (33) to be valid. This constraint does not impose a significant restriction on the selection of the pair of $x_{0},y_{0}$, as for a small Tx array and a large Rx array, the lower limit for the distance $d_{0}$ is relatively small. Furthermore, this constraint contributes to the desired value of the ratio $[(\rho-\rho’)/(\tilde{\rho}-\tilde{\rho}’)l_{T}]$, i.e., the greater the distance $d_{0}$, the greater the ratio becomes. The second observation is that the larger the values of $L_{T}$ and $L_{R}$, the larger the number of DoF, as expected and also shown in Fig. 3(c) and 3(d). However, the value of $L_{T}$ also determines the effective length $l_{T}$ that is desirable to keep it at relatively low values. If $L_{T}\ll d_{0}$, the ratio $[(\rho-\rho’)/(\tilde{\rho}-\tilde{\rho}’)l_{T}]$ is kept large and achieves the orthogonality of the focusing functions. The third observation is that the rotation angles, $\theta_{T}$ and $\theta_{R}$, not only determine the visibility of the two arrays but also alter the projected lengths of the two linear arrays onto the perpendicular line connecting the centers of the two arrays. This has a direct consequence on the number of DoF as also observed by comparing Figs. 3(a) and 3(b). The partial visibility of the Tx or Rx arrays directly affects the number of DoF in the specific geometric setup. In fact, the partial visibility of the Tx array implies a reduced length compared to $L_{T}$, which in turn changes the position of the zeros of the kernel on the Rx array. Furthermore, the number of DoF increases with $l_{T}$ according to (38). The same applies to the reduced effective length of the Rx array, which affects the range of the $\zeta$ variable and, consequently, the number of DoF. Using the LoS channel model and performing a singular value decomposition (SVD) to the channel matrix, one can obtain the DoF for a specific geometric setup of the two arrays. The channel matrix is a non-square matrix that contains the values of the scalar free space Green’s function that models the response at a point, $\zeta_{m}$, on the Rx array from a wave transmitted from a point $\eta_{n}$ on the Tx array. Specifically, using the retarded solution to the wave equation, i.e., the outgoing wave, one may write $G(\zeta_{m},\eta_{n})=\frac{1}{4\pi}\frac{\exp{(-jkr_{mn}})}{r_{mn}}$, where $r_{mn}$ is given by (1) for $\zeta=\zeta_{m}$ and $\eta=\eta_{n}$. Therefore, the distance from a Tx point to an Rx point is calculated without any approximation. Fig. 5 presents the eigenvalues normalized to the maximum eigenvalue versus their index, for the set of parameters used in Fig. 4, and for three values of the angle $\theta_{R}$. It is evident that there is a threshold for the number of DoF after which the eigenvalues decrease rapidly. The precise number of DoF that corresponds to the knee of the curves depends on the power coupling strength of the modes. The sum rule on the eigenvalues provides the total power strength of the eigenvalues and by setting a percentage of the sum rule to be guaranteed, one can determine the number of orthogonal modes to retain [52]. Orthogonality is directly related to the zeros of the kernel function, through the inner products in the Hilbert spaces. The eigenvalues of the modes after the knee indicate a weak strength and the inability to exploit additional DoF. Consequently, the knee corresponds to the effective “diffraction” limit, or the so-called effective number of DoF, and its value almost coincides with the value displayed in Fig. 4 for the corresponding values of $\theta_{R}$ and for a threshold level of 99% of the sum rule. The 11th eigenvalue for $\theta_{R}=53^{o}$ is relatively weak, $|s_{11}|^{2}=0.194$, compared to the 10th, $|s_{10}|^{2}=0.416$, where $|s_{j}|$ denotes the magnitude of the $j$-th singular value. The percentage of cumulative strength of the first ten eigenvalues in approximately 96%.
@@ IV-A Insights on the Calculation of the Number of DoF
Figure 5: The DoF for $\theta_{T}=\pi/2$, $x_{0}=-5$ m, $y_{0}=5$ m, $L_{T}=0.2$ m, $L_{R}=5$ m, $f=30\,$GHz as calculated using the SVD of the LoS channel.
IV-B A Geometric Setup of Interest
In this subsection, a practical geometric setup is considered where a small array is located in front of a large array, which resembles the placement of user equipment in front of a radio unit. This case study is of interest since, for a range of values of the variables $x_{0}$ and $\theta_{T}$, it provides a paraxial setup, i.e., the midpoints of the two arrays are perfectly aligned along the same line and thus $y_{0}=0$, whereas for other values of $x_{0}$ and $\theta_{T}$ the center of the Rx is shifted, leading to a non-paraxial case. This is due to the partial visibility of the Rx array and thus the need to calculate the DoF with respect to the new center. This setup will also be utilized in the next section to study the statistical behavior of the DoF. Moreover, assume that $x_{0}\geq 1.2(L_{T}+L_{R})$, that is, Rx is in the right half-plane, and $\theta_{R}=\pi$. The coordinate $x_{0}$ and the rotation angle of the TX surface determine the visibility condition, and consequently the number of DoF between the two surfaces. Therefore, there are four possible cases: i) the Rx array is not visible from the Tx and the DoF is 0, ii) the Rx array is partially visible from the Tx array and the endpoint $R^{+}$ is visible, iii) the Rx array is fully visible from the Tx, and iv) the Rx array is partially visible from the Tx and the endpoint $R^{-}$ is visible. The first case occurs for $-\pi<\theta_{T}\leq a_{+}^{max}-\frac{\pi}{2}$ and for $a_{-}^{max}+\frac{\pi}{2}\leq\theta_{T}\leq\pi$, where $a_{+}^{max}=\arctan\gamma_{+}=\arctan(-L_{R}/2x_{0})$ and $a_{-}^{max}=\arctan\gamma_{-}=\arctan(L_{R}/2x_{0})=-a_{+}^{max}$. Then,
@@ IV-B A Geometric Setup of Interest
$m$ $=1+\frac{L_{T}}{\lambda}[\sin\left(\theta_{T}-a_{+}\right)-\sin\left(\theta_{T}-a_{-}\right)]$ (47)
$\mathop{=}\limits^{a_{+}=-a_{-}}1+\frac{2L_{T}}{\lambda}\cos\theta_{T}\sin a_{-}.$
@@ V-A Stochastic Geometry Framework
$m=1+\frac{L_{T}}{\lambda}|\sin(\theta_{T}-a_{+})-\sin(\theta_{T}-a_{-})|=1+\frac{L_{T}}{\lambda}|\rho_{+}-\rho_{-}|.$ (49)
@@ V-C1 Partial Visibility of the Rx Array
for $m\in\Big{[}0,2C_{L_{T},\lambda}\frac{L_{R}^{2}}{L_{R}^{2}+4x_{0}^{2}}\Big{]}$. Proof. Assume that $R^{+}$ is visible. Conditioned on $x_{0}$, one can set $f_{x_{0}}(x_{0})=\delta(x_{0})$ in (50), which provides the PDF $f_{\rho_{-}}(\rho)$. The limits of the range for $m$ are obtained with the aid of (60) shown in Appendix A. Therefore, $f_{m|x_{0},R^{+}}(m)=f_{m|R^{+}}(m)\Big{|}_{f_{x_{0}}(x_{0})=\delta(x_{0})}$ and similarly $f_{m|x_{0},R^{-}}(m)=f_{m|R^{-}}(m)\Big{|}_{f_{x_{0}}(x_{0})=\delta(x_{0})}$. Finally, $f_{m|x_{0},R^{+}}(m)$ and $f_{m|x_{0},R^{i}}(m)$ have the same statistical behavior. $\blacksquare$
@@ V-C2 Full Visibility of the Rx Array
Conditioned on $x_{0}$, and with full visibility of the Rx array, the conditional PDF $f_{m|x_{0},FV}(m)$ of the number of DoF is given by Lemma 3 as $f_{m|x_{0},FV}(m)=f_{m|x_{0}}(m)$. Conditioned on $x_{0}$, the PDF of the number of DoF is given by the following theorem. Theorem 4: Conditioned on $x_{0}$, the conditional PDF $f_{m|x_{0}}(m)$ of the number of DoF, $m$, is given by (58), shown at the bottom of the next page, where $\mathcal{V}_{R^{+}}=\mathcal{V}_{R^{-}}=\frac{\arctan(L_{R}/2x_{0})}{\pi}$ denotes the probability that $R^{+/-}$ is visible and $\mathcal{V}_{FV}=\frac{\pi-2\arctan(L_{R}/2x_{0})}{2\pi}$ denotes the probability that full visibility is achievable.
@@
VI Numerical Results and Discussion
VI-A Deterministic Analysis
Because there are seven variables involved in the calculation of the number of DoF, namely $\theta_{T},\theta_{R},x_{0},y_{0},L_{T},L_{R},f$, the number of possible combinations is extremely large. Thus, in this subsection, representative cases are examined and numerical results are given. In all results presented, the frequency is set to $f=30\,$GHz, and the length of the transmitting surface is set to $L_{T}=0.2$ meters. In Fig. 7, the number of DoF for three different Rx locations is given assuming that a large receiving array of length $L_{R}=5$ meters is used. The number of DoF is given versus $\theta_{R}\in(-\pi/2,\pi/2]$, for various values of $\theta_{T}$. The line that connects the centers of the two arrays is the $x$-axis in the first plot, the $y$-axis in the second plot, and it forms an angle of $\pi/4$ with the $x$-axis in the third plot. All the depicted values match the number of zeros of the kernel function as shown in Fig. 3. The first observation is that a different maximum number of DoF is obtained for different values of $\theta_{T}$. Notably, the maximum value of the number of DoF for $\theta_{T}=0^{o}$ in Fig. 7(a) and $\theta_{T}=90^{o}$ in 7(b), matches the heuristic number of DoF in th eparaxial setting, i.e., $m\approx\frac{L_{T}L_{R}}{\lambda d_{0}}$ [52]. A second observation is that when the distance between the two arrays is reduced from $d_{0}=10\,m$ in Fig. 7(a) to $d_{0}\approx 7\,m$ in Fig. 7(c), the maximum number of DoF increases. Another observation is the high number of modes available in the considered settings, for a small transmitting array and for a wide range of angles $\theta_{T}$ and $\theta_{R}$. An interesting conclusion based on the results obtained is that, in a multi-user scenario where the users are equipped with small arrays and the BS with a large one, the knowledge of the location and the relative orientation of the arrays may be used to design efficient resource allocation strategies, based on the number of DoF attainable at each small array.
A comparison of the number of DoF calculated by the proposed method and the SVD is shown in Fig. 8. A round to the nearest integer has been applied to the kernel-based method for comparison with the SVD. The sub-figures correspond to two indicative values of $\theta_{T}$. The SVD was applied considering the same visibility conditions and effective length as calculated by Algorithm 1, and a threshold value 96% of the sum-rule. The 11-th mode shown in Fig. 8(a) and 8(b) corresponds to a relatively weak eigenvalue.